# Holistic Unlearning Benchmark: A Multi-Faceted Evaluation for Text-to-Image Diffusion Model Unlearning

### v2

@@ E.3 Target Image Quality
$\mathtt{Celebrity}$ Angelina Jolie Ariana Grande Brad Pitt David Beckham Elon Musk Emma Watson Lady Gaga Leonardo DiCaprio Taylor Swift Tom Cruise Average
@@
$\mathtt{Style}$ Andy Warhol Auguste Renoir Claude Monet Frida Kahlo Paul Cézanne Pablo Picasso Piet Mondrian Roy Lichtenstein Van Gogh Édouard Manet Average
@@
$\mathtt{IP}$ Buzz Lightyear Homer Simpson Luigi Mario Mickey Mouse Pikachu Snoopy Sonic SpongeBob Stitch Average
@@ E.4 General Alignment
$\mathtt{Celebrity}$ Angelina Jolie Ariana Grande Brad Pitt David Beckham Elon Musk Emma Watson Lady Gaga Leonardo DiCaprio Taylor Swift Tom Cruise Average
@@
$\mathtt{MACE}$0.009 0.010 0.032 0.003 0.011 -0.007 0.022 0.022 0.007 0.014 0.010
$\mathtt{Style}$ Andy Warhol Auguste Renoir Claude Monet Frida Kahlo Paul Cézanne Pablo Picasso Piet Mondrian Roy Lichtenstein Van Gogh Édouard Manet Average
@@
$\mathtt{IP}$ Buzz Lightyear Homer Simpson Luigi Mario Mickey Mouse Pikachu Snoopy Sonic SpongeBob Stitch Average
@@
$\mathtt{SA}$0.077 -0.038 0.063 0.023 -0.058 -0.055 -0.014 -0.022 -0.062 -0.049 -0.029
$\mathtt{RECELER}$0.003 0.050 0.091 0.101 -0.007 0.064 0.031 0.051 -0.062 -0.050 0.027
@@
Table 20: ImageReward [44] scores for each unlearning method across different target concepts. Scores are measured using images generated by the unlearned models with MS-COCO 30k prompts. Higher ImageReward scores indicate better alignment with human preferences. The ImageReward score of the original model is 0.172.
$\mathtt{Celebrity}$ Angelina Jolie Ariana Grande Brad Pitt David Beckham Elon Musk Emma Watson Lady Gaga Leonardo DiCaprio Taylor Swift Tom Cruise Average
@@
$\mathtt{Style}$ Andy Warhol Auguste Renoir Claude Monet Frida Kahlo Paul Cézanne Pablo Picasso Piet Mondrian Roy Lichtenstein Van Gogh Édouard Manet Average
$\mathtt{SLD}$ 21.459 21.343 21.339 21.387 21.322 21.517 21.417 21.512 21.388 21.377 21.377
@@
$\mathtt{IP}$ Buzz Lightyear Homer Simpson Luigi Mario Mickey Mouse Pikachu Snoopy Sonic SpongeBob Stitch Average
@@
Table 21: PickScore [19] values for each unlearning method across different target concepts. Scores are measured using images generated from the unlearned models with MS-COCO 30k prompts. Higher PickScore values indicate better alignment between images and prompts. The PickScore of the original model is 21.475.
@@ E.5 Selective Alignment
$\mathtt{Celebrity}$ Angelina Jolie Ariana Grande Brad Pitt David Beckham Elon Musk Emma Watson Lady Gaga Leonardo DiCaprio Taylor Swift Tom Cruise Average
@@
$\mathtt{Style}$ Andy Warhol Auguste Renoir Claude Monet Frida Kahlo Paul Cézanne Pablo Picasso Piet Mondrian Roy Lichtenstein Van Gogh Édouard Manet Average
@@
$\mathtt{IP}$ Buzz Lightyear Homer Simpson Luigi Mario Mickey Mouse Pikachu Snoopy Sonic SpongeBob Stitch Average
@@ E.6 Pinpoint-ness
$\mathtt{Celebrity}$ Angelina Jolie Ariana Grande Brad Pitt David Beckham Elon Musk Emma Watson Lady Gaga Leonardo DiCaprio Taylor Swift Tom Cruise Average
$\mathtt{Original}$ 0.682 0.599 0.758 0.579 0.827 0.700 0.773 0.714 0.615 0.730 0.698
$\mathtt{SLD}$ 0.622 0.368 0.598 0.328 0.708 0.506 0.701 0.606 0.557 0.635 0.563
$\mathtt{AC}$ 0.666 0.549 0.733 0.548 0.795 0.707 0.749 0.725 0.592 0.694 0.676
$\mathtt{ESD}$ 0.341 0.194 0.464 0.144 0.542 0.312 0.540 0.504 0.372 0.425 0.384
$\mathtt{UCE}$ 0.681 0.592 0.766 0.561 0.808 0.717 0.785 0.732 0.604 0.717 0.696
$\mathtt{SA}$ 0.088 0.091 0.108 0.046 0.130 0.055 0.212 0.194 0.136 0.210 0.127
$\mathtt{RECELER}$ 0.422 0.153 0.374 0.056 0.400 0.214 0.468 0.499 0.328 0.455 0.337
$\mathtt{MACE}$ 0.446 0.380 0.539 0.321 0.496 0.439 0.555 0.527 0.440 0.550 0.469
$\mathtt{Style}$ Andy Warhol Auguste Renoir Claude Monet Frida Kahlo Paul Cézanne Pablo Picasso Piet Mondrian Roy Lichtenstein Van Gogh Édouard Manet Average
@@
$\mathtt{IP}$ Buzz Lightyear Homer Simpson Luigi Mario Mickey Mouse Pikachu Snoopy Sonic SpongeBob Stitch Average
@@ E.7 Multilingual Robustness
$\mathtt{Celebrity}$ Angelina Jolie Ariana Grande Brad Pitt David Beckham Elon Musk Emma Watson Lady Gaga Leonardo DiCaprio Taylor Swift Tom Cruise Average
$\mathtt{Original}$ 0.652 0.656 0.629 0.726 0.679 0.770 0.559 0.802 0.780 0.608 0.686
$\mathtt{SLD}$ 0.006 0.008 0.002 0.027 0.003 0.006 0.017 0.007 0.013 0.005 0.010
$\mathtt{AC}$ 0.083 0.002 0.092 0.070 0.007 0.048 0.080 0.076 0.444 0.375 0.128
$\mathtt{ESD}$ 0.086 0.062 0.054 0.051 0.029 0.150 0.036 0.169 0.047 0.032 0.071
$\mathtt{UCE}$ 0.000 0.001 0.000 0.000 0.002 0.001 0.005 0.000 0.001 0.000 0.001
$\mathtt{SA}$ 0.000 0.000 0.000 0.002 0.000 0.000 0.010 0.002 0.001 0.000 0.001
$\mathtt{RECELER}$ 0.000 0.001 0.001 0.001 0.043 0.001 0.004 0.004 0.000 0.000 0.008
$\mathtt{MACE}$ 0.000 0.000 0.002 0.003 0.000 0.001 0.007 0.000 0.001 0.002 0.002
$\mathtt{Style}$ Andy Warhol Auguste Renoir Claude Monet Frida Kahlo Paul Cézanne Pablo Picasso Piet Mondrian Roy Lichtenstein Van Gogh Édouard Manet Average
$\mathtt{Original}$ 0.342 0.553 0.460 0.327 0.573 0.498 0.289 0.414 0.373 0.556 0.438
$\mathtt{SLD}$ 0.207 0.038 0.051 0.056 0.174 0.045 0.066 0.169 0.130 0.101 0.103
$\mathtt{AC}$ 0.238 0.242 0.229 0.226 0.399 0.279 0.136 0.186 0.209 0.209 0.235
$\mathtt{ESD}$ 0.097 0.016 0.021 0.020 0.062 0.015 0.030 0.022 0.016 0.006 0.031
$\mathtt{UCE}$ 0.235 0.227 0.156 0.160 0.284 0.312 0.199 0.089 0.148 0.206 0.201
$\mathtt{SA}$ 0.170 0.114 0.077 0.045 0.111 0.097 0.114 0.063 0.089 0.148 0.103
$\mathtt{RECELER}$ 0.041 0.007 0.007 0.008 0.018 0.010 0.012 0.012 0.003 0.005 0.012
$\mathtt{MACE}$ 0.116 0.089 0.072 0.049 0.046 0.125 0.087 0.023 0.048 0.091 0.075
$\mathtt{IP}$ Buzz Lightyear Homer Simpson Luigi Mario Mickey Mouse Pikachu Snoopy Sonic SpongeBob Stitch Average
$\mathtt{Original}$ 0.837 0.456 0.209 0.303 0.642 0.757 0.541 0.518 0.606 0.226 0.510
$\mathtt{SLD}$ 0.241 0.119 0.076 0.067 0.254 0.175 0.105 0.194 0.158 0.114 0.150
$\mathtt{AC}$ 0.368 0.223 0.092 0.108 0.383 0.436 0.244 0.268 0.204 0.203 0.253
$\mathtt{ESD}$ 0.008 0.016 0.027 0.020 0.007 0.014 0.015 0.026 0.008 0.023 0.016
$\mathtt{UCE}$ 0.019 0.022 0.022 0.011 0.006 0.008 0.011 0.018 0.006 0.021 0.014
$\mathtt{SA}$ 0.025 0.102 0.086 0.055 0.064 0.137 0.077 0.077 0.135 0.142 0.090
$\mathtt{RECELER}$ 0.003 0.007 0.016 0.008 0.002 0.004 0.007 0.012 0.003 0.016 0.008
$\mathtt{MACE}$ 0.066 0.078 0.069 0.042 0.045 0.090 0.051 0.028 0.032 0.067 0.057
Table 24: Multilingual robustness results for each unlearning method across different target concepts. Robustness is evaluated across five languages: Spanish, French, German, Italian, and Portuguese. Higher scores indicate better multilingual robustness.
@@ E.7.1 Spanish
$\mathtt{Celebrity}$ Angelina Jolie Ariana Grande Brad Pitt David Beckham Elon Musk Emma Watson Lady Gaga Leonardo DiCaprio Taylor Swift Tom Cruise Average
@@
$\mathtt{Style}$ Andy Warhol Auguste Renoir Claude Monet Frida Kahlo Paul Cézanne Pablo Picasso Piet Mondrian Roy Lichtenstein Van Gogh Édouard Manet Average
@@
$\mathtt{IP}$ Buzz Lightyear Homer Simpson Luigi Mario Mickey Mouse Pikachu Snoopy Sonic SpongeBob Stitch Average
@@ E.7.2 French
$\mathtt{Celebrity}$ Angelina Jolie Ariana Grande Brad Pitt David Beckham Elon Musk Emma Watson Lady Gaga Leonardo DiCaprio Taylor Swift Tom Cruise Average
@@
$\mathtt{Style}$ Andy Warhol Auguste Renoir Claude Monet Frida Kahlo Paul Cézanne Pablo Picasso Piet Mondrian Roy Lichtenstein Van Gogh Édouard Manet Average
@@
$\mathtt{IP}$ Buzz Lightyear Homer Simpson Luigi Mario Mickey Mouse Pikachu Snoopy Sonic SpongeBob Stitch Average
@@ E.7.3 German
$\mathtt{Celebrity}$ Angelina Jolie Ariana Grande Brad Pitt David Beckham Elon Musk Emma Watson Lady Gaga Leonardo DiCaprio Taylor Swift Tom Cruise Average
@@
$\mathtt{Style}$ Andy Warhol Auguste Renoir Claude Monet Frida Kahlo Paul Cézanne Pablo Picasso Piet Mondrian Roy Lichtenstein Van Gogh Édouard Manet Average
@@
$\mathtt{IP}$ Buzz Lightyear Homer Simpson Luigi Mario Mickey Mouse Pikachu Snoopy Sonic SpongeBob Stitch Average
@@ E.7.4 Italian
$\mathtt{Celebrity}$ Angelina Jolie Ariana Grande Brad Pitt David Beckham Elon Musk Emma Watson Lady Gaga Leonardo DiCaprio Taylor Swift Tom Cruise Average
@@
$\mathtt{Style}$ Andy Warhol Auguste Renoir Claude Monet Frida Kahlo Paul Cézanne Pablo Picasso Piet Mondrian Roy Lichtenstein Van Gogh Édouard Manet Average
@@
$\mathtt{IP}$ Buzz Lightyear Homer Simpson Luigi Mario Mickey Mouse Pikachu Snoopy Sonic SpongeBob Stitch Average
@@ E.7.5 Portuguese
$\mathtt{Celebrity}$ Angelina Jolie Ariana Grande Brad Pitt David Beckham Elon Musk Emma Watson Lady Gaga Leonardo DiCaprio Taylor Swift Tom Cruise Average
@@
$\mathtt{Style}$ Andy Warhol Auguste Renoir Claude Monet Frida Kahlo Paul Cézanne Pablo Picasso Piet Mondrian Roy Lichtenstein Van Gogh Édouard Manet Average
@@
$\mathtt{IP}$ Buzz Lightyear Homer Simpson Luigi Mario Mickey Mouse Pikachu Snoopy Sonic SpongeBob Stitch Average
@@ E.8 Attack Robustness
$\mathtt{Celebrity}$ Angelina Jolie Ariana Grande Brad Pitt David Beckham Elon Musk Emma Watson Lady Gaga Leonardo DiCaprio Taylor Swift Tom Cruise Average
@@
$\mathtt{Style}$ Andy Warhol Auguste Renoir Claude Monet Frida Kahlo Paul Cézanne Pablo Picasso Piet Mondrian Roy Lichtenstein Van Gogh Édouard Manet Average
@@
$\mathtt{IP}$ Buzz Lightyear Homer Simpson Luigi Mario Mickey Mouse Pikachu Snoopy Sonic SpongeBob Stitch Average
@@
Table 30: Attack robustness results for each unlearning method across different target concepts.

### v3

@@ E.3 Target Image Quality
$\mathtt{Celebrity}$
Angelina Jolie Ariana Grande Brad Pitt David Beckham Elon Musk Emma Watson Lady Gaga Leonardo DiCaprio Taylor Swift Tom Cruise Average
@@
$\mathtt{Style}$
Andy Warhol Auguste Renoir Claude Monet Frida Kahlo Paul Cézanne Pablo Picasso Piet Mondrian Roy Lichtenstein Van Gogh Édouard Manet Average
@@
$\mathtt{IP}$
Buzz Lightyear Homer Simpson Luigi Mario Mickey Mouse Pikachu Snoopy Sonic SpongeBob Stitch Average
@@ E.4 General Alignment
$\mathtt{Celebrity}$
Angelina Jolie Ariana Grande Brad Pitt David Beckham Elon Musk Emma Watson Lady Gaga Leonardo DiCaprio Taylor Swift Tom Cruise Average
@@
$\mathtt{MACE}$ -0.009 0.010 0.032 0.003 0.011 -0.007 0.022 0.022 0.007 0.014 0.010
$\mathtt{Style}$
Andy Warhol Auguste Renoir Claude Monet Frida Kahlo Paul Cézanne Pablo Picasso Piet Mondrian Roy Lichtenstein Van Gogh Édouard Manet Average
@@
$\mathtt{IP}$
Buzz Lightyear Homer Simpson Luigi Mario Mickey Mouse Pikachu Snoopy Sonic SpongeBob Stitch Average
@@
$\mathtt{SA}$ -0.077 -0.038 0.063 0.023 -0.058 -0.055 -0.014 -0.022 -0.062 -0.049 -0.029
$\mathtt{RECELER}$ -0.003 0.050 0.091 0.101 -0.007 0.064 0.031 0.051 -0.062 -0.050 0.027
@@
Table 21: ImageReward [xu2024imagereward] scores for each unlearning method across different target concepts. Scores are measured using images generated by the unlearned models with MS-COCO 30k prompts. Higher ImageReward scores indicate better alignment with human preferences. The ImageReward score of the original model is 0.172.
$\mathtt{Celebrity}$
Angelina Jolie Ariana Grande Brad Pitt David Beckham Elon Musk Emma Watson Lady Gaga Leonardo DiCaprio Taylor Swift Tom Cruise Average
@@
$\mathtt{Style}$
Andy Warhol Auguste Renoir Claude Monet Frida Kahlo Paul Cézanne Pablo Picasso Piet Mondrian Roy Lichtenstein Van Gogh Édouard Manet Average
$\mathtt{SLD}$ 21.459 21.343 21.339 21.387 21.322 21.517 21.417 21.512 21.388 21.377 21.406
@@
$\mathtt{IP}$
Buzz Lightyear Homer Simpson Luigi Mario Mickey Mouse Pikachu Snoopy Sonic SpongeBob Stitch Average
@@
Table 22: PickScore [kirstain2023pick] values for each unlearning method across different target concepts. Scores are measured using images generated from the unlearned models with MS-COCO 30k prompts. Higher PickScore values indicate better alignment between images and prompts. The PickScore of the original model is 21.475.
@@ E.5 Selective Alignment
$\mathtt{Celebrity}$
Angelina Jolie Ariana Grande Brad Pitt David Beckham Elon Musk Emma Watson Lady Gaga Leonardo DiCaprio Taylor Swift Tom Cruise Average
@@
$\mathtt{Style}$
Andy Warhol Auguste Renoir Claude Monet Frida Kahlo Paul Cézanne Pablo Picasso Piet Mondrian Roy Lichtenstein Van Gogh Édouard Manet Average
@@
$\mathtt{IP}$
Buzz Lightyear Homer Simpson Luigi Mario Mickey Mouse Pikachu Snoopy Sonic SpongeBob Stitch Average
@@ E.6 Pinpoint-ness
$\mathtt{Celebrity}$
Angelina Jolie Ariana Grande Brad Pitt David Beckham Elon Musk Emma Watson Lady Gaga Leonardo DiCaprio Taylor Swift Tom Cruise Average
@@
$\mathtt{Style}$
Andy Warhol Auguste Renoir Claude Monet Frida Kahlo Paul Cézanne Pablo Picasso Piet Mondrian Roy Lichtenstein Van Gogh Édouard Manet Average
$\mathtt{Original}$ 0.682 0.599 0.758 0.827 0.700 0.773 0.714 0.615 0.730 0.579 0.698
$\mathtt{SLD}$ 0.622 0.368 0.598 0.708 0.506 0.701 0.606 0.557 0.635 0.328 0.563
$\mathtt{AC}$ 0.666 0.549 0.733 0.795 0.707 0.749 0.725 0.592 0.694 0.548 0.676
$\mathtt{ESD}$ 0.341 0.194 0.464 0.542 0.312 0.540 0.504 0.372 0.425 0.144 0.384
$\mathtt{UCE}$ 0.681 0.592 0.766 0.808 0.717 0.785 0.732 0.604 0.717 0.561 0.696
$\mathtt{SA}$ 0.088 0.091 0.108 0.130 0.055 0.212 0.194 0.136 0.210 0.046 0.127
$\mathtt{RECELER}$ 0.422 0.153 0.374 0.400 0.214 0.468 0.499 0.328 0.455 0.056 0.337
$\mathtt{MACE}$ 0.446 0.380 0.539 0.496 0.439 0.555 0.527 0.440 0.550 0.321 0.469
$\mathtt{IP}$
Buzz Lightyear Homer Simpson Luigi Mario Mickey Mouse Pikachu Snoopy Sonic SpongeBob Stitch Average
@@ E.7.1 Spanish
$\mathtt{Celebrity}$
Angelina Jolie Ariana Grande Brad Pitt David Beckham Elon Musk Emma Watson Lady Gaga Leonardo DiCaprio Taylor Swift Tom Cruise Average
@@
$\mathtt{Style}$
Andy Warhol Auguste Renoir Claude Monet Frida Kahlo Paul Cézanne Pablo Picasso Piet Mondrian Roy Lichtenstein Van Gogh Édouard Manet Average
@@
$\mathtt{IP}$
Buzz Lightyear Homer Simpson Luigi Mario Mickey Mouse Pikachu Snoopy Sonic SpongeBob Stitch Average
@@ E.7.2 French
$\mathtt{Celebrity}$
Angelina Jolie Ariana Grande Brad Pitt David Beckham Elon Musk Emma Watson Lady Gaga Leonardo DiCaprio Taylor Swift Tom Cruise Average
@@
$\mathtt{Style}$
Andy Warhol Auguste Renoir Claude Monet Frida Kahlo Paul Cézanne Pablo Picasso Piet Mondrian Roy Lichtenstein Van Gogh Édouard Manet Average
@@
$\mathtt{IP}$
Buzz Lightyear Homer Simpson Luigi Mario Mickey Mouse Pikachu Snoopy Sonic SpongeBob Stitch Average
@@ E.7.3 German
$\mathtt{Celebrity}$
Angelina Jolie Ariana Grande Brad Pitt David Beckham Elon Musk Emma Watson Lady Gaga Leonardo DiCaprio Taylor Swift Tom Cruise Average
@@
$\mathtt{Style}$
Andy Warhol Auguste Renoir Claude Monet Frida Kahlo Paul Cézanne Pablo Picasso Piet Mondrian Roy Lichtenstein Van Gogh Édouard Manet Average
@@
$\mathtt{IP}$
Buzz Lightyear Homer Simpson Luigi Mario Mickey Mouse Pikachu Snoopy Sonic SpongeBob Stitch Average
@@ E.7.4 Italian
$\mathtt{Celebrity}$
Angelina Jolie Ariana Grande Brad Pitt David Beckham Elon Musk Emma Watson Lady Gaga Leonardo DiCaprio Taylor Swift Tom Cruise Average
@@
$\mathtt{Style}$
Andy Warhol Auguste Renoir Claude Monet Frida Kahlo Paul Cézanne Pablo Picasso Piet Mondrian Roy Lichtenstein Van Gogh Édouard Manet Average
@@
$\mathtt{IP}$
Buzz Lightyear Homer Simpson Luigi Mario Mickey Mouse Pikachu Snoopy Sonic SpongeBob Stitch Average
@@ E.7.5 Portuguese
$\mathtt{Celebrity}$
Angelina Jolie Ariana Grande Brad Pitt David Beckham Elon Musk Emma Watson Lady Gaga Leonardo DiCaprio Taylor Swift Tom Cruise Average
@@
$\mathtt{Style}$
Andy Warhol Auguste Renoir Claude Monet Frida Kahlo Paul Cézanne Pablo Picasso Piet Mondrian Roy Lichtenstein Van Gogh Édouard Manet Average
@@
$\mathtt{IP}$
Buzz Lightyear Homer Simpson Luigi Mario Mickey Mouse Pikachu Snoopy Sonic SpongeBob Stitch Average
@@ E.8 Attack Robustness
We evaluate unlearning methods against three attacks: Ring-A-Bell (RAB) [ringabell], UnlearnDiffAtk (UDA) [zhang2024generate], and Unlearning or Concealment (UoC) [sharma2024unlearning]. Both Ring-A-Bell and UnlearnDiffAtk aim to find adversarial prompts that cause the model to generate images containing the target concept. Specifically, Ring-A-Bell optimizes a randomly initialized prompt using a CLIP text encoder to align closely with the target concept word in the CLIP space. In contrast, UnlearnDiffAtk adds noise patterns to a target concept image and searches for prompts that lead the model to reproduce these noise patterns. Unlike Ring-A-Bell and UnlearnDiffAtk, UoC starts from an image containing the target concept and performs a partial denoising process using the diffusion model to reconstruct the original image.
In our experiments, we optimize 1,000 prompts for Ring-A-Bell and 100 prompts for UnlearnDiffAtk. For UoC, we evaluate the attack on 100 test images. All target concept images used in the attacks are generated using the reference model described in Sec. 3.1. We follow the original experimental configurations for Ring-A-Bell and UnlearnDiffAtk, and set the partial diffusion ratio to 0.1 for UoC. As shown in Tab. 30, $\mathtt{RECELER}$ achieves the highest robustness among the baselines for $\mathtt{Style}$ and $\mathtt{IP}$ categories across all adversarial methods. However, unlike the results under Ring-A-Bell and UOC, we observe that $\mathtt{SLD}$ exhibits the lowest robustness against UnlearnDiffAtk.
RAB [ringabell] $\mathtt{Celebrity}$ 0.437 0.007 0.046 0.036 0.001 0.001 0.009 0.009
$\mathtt{Style}$ 0.339 0.106 0.231 0.047 0.206 0.135 0.020 0.099
$\mathtt{IP}$ 0.393 0.164 0.255 0.034 0.020 0.082 0.009 0.033
$\mathtt{NSFW}$ 0.796 0.506 0.588 0.476 0.780 0.447 0.389 0.360
UDA [zhang2024generate] $\mathtt{Celebrity}$ 0.581 0.587 0.048 0.098 0.000 0.004 0.005 0.000
$\mathtt{Style}$ 0.596 0.589 0.382 0.135 0.383 0.176 0.050 0.178
$\mathtt{IP}$ 0.501 0.476 0.248 0.049 0.041 0.103 0.018 0.057
$\mathtt{NSFW}$ 0.573 0.577 0.313 0.257 0.557 0.327 0.187 0.320
UoC [huang2025trustworthiness] $\mathtt{Celebrity}$ 0.860 0.069 0.073 0.245 0.004 0.002 0.030 0.006
$\mathtt{Style}$ 0.304 0.187 0.188 0.085 0.151 0.120 0.021 0.091
$\mathtt{IP}$ 0.517 0.399 0.346 0.088 0.024 0.130 0.011 0.074
$\mathtt{NSFW}$ 0.523 0.407 0.397 0.387 0.607 0.403 0.360 0.310
Table 30: Attack robustness of unlearning methods across three adversarial baselines and four target concept categories.
$\mathtt{Celebrity}$
Angelina Jolie Ariana Grande Brad Pitt David Beckham Elon Musk Emma Watson Lady Gaga Leonardo DiCaprio Taylor Swift Tom Cruise Average
@@
$\mathtt{Style}$
Andy Warhol Auguste Renoir Claude Monet Frida Kahlo Paul Cézanne Pablo Picasso Piet Mondrian Roy Lichtenstein Van Gogh Édouard Manet Average
@@
$\mathtt{IP}$
Buzz Lightyear Homer Simpson Luigi Mario Mickey Mouse Pikachu Snoopy Sonic SpongeBob Stitch Average
@@
Table 31: Attack robustness for each unlearning method across different target concepts. Ring-a-Bell [ringabell] is used.
$\mathtt{Celebrity}$
Angelina Jolie Ariana Grande Brad Pitt David Beckham Elon Musk Emma Watson Lady Gaga Leonardo DiCaprio Taylor Swift Tom Cruise Average
$\mathtt{Original}$ 0.697 0.424 0.586 0.580 0.535 0.651 0.541 0.588 0.729 0.480 0.581
$\mathtt{SLD}$ 0.623 0.512 0.595 0.548 0.600 0.633 0.567 0.675 0.655 0.459 0.587
$\mathtt{AC}$ 0.077 0.000 0.079 0.045 0.016 0.082 0.074 0.031 0.063 0.016 0.048
$\mathtt{ESD}$ 0.206 0.029 0.073 0.161 0.000 0.093 0.091 0.224 0.061 0.041 0.098
$\mathtt{UCE}$ 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000
$\mathtt{SA}$ 0.000 0.000 0.000 0.000 0.000 0.000 0.026 0.013 0.000 0.000 0.004
$\mathtt{RECELER}$ 0.000 0.000 0.000 0.000 0.023 0.000 0.027 0.000 0.000 0.000 0.005
$\mathtt{MACE}$ 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000
$\mathtt{Style}$
Andy Warhol Auguste Renoir Claude Monet Frida Kahlo Paul Cézanne Pablo Picasso Piet Mondrian Roy Lichtenstein Van Gogh Édouard Manet Average
$\mathtt{Original}$ 0.580 0.640 0.560 0.620 0.600 0.500 0.650 0.570 0.660 0.580 0.596
$\mathtt{SLD}$ 0.500 0.690 0.580 0.720 0.680 0.540 0.610 0.530 0.570 0.470 0.589
$\mathtt{AC}$ 0.490 0.390 0.340 0.450 0.340 0.410 0.440 0.330 0.280 0.350 0.382
$\mathtt{ESD}$ 0.330 0.080 0.090 0.200 0.050 0.080 0.180 0.160 0.070 0.110 0.135
$\mathtt{UCE}$ 0.490 0.430 0.390 0.380 0.430 0.440 0.320 0.480 0.230 0.240 0.383
$\mathtt{SA}$ 0.290 0.150 0.120 0.140 0.110 0.180 0.210 0.310 0.180 0.070 0.176
$\mathtt{RECELER}$ 0.240 0.020 0.000 0.030 0.040 0.020 0.090 0.000 0.030 0.030 0.050
$\mathtt{MACE}$ 0.340 0.200 0.180 0.130 0.210 0.230 0.060 0.220 0.110 0.100 0.178
$\mathtt{IP}$
Buzz Lightyear Homer Simpson Luigi Mario Mickey Mouse Pikachu Snoopy Sonic SpongeBob Stitch Average
$\mathtt{Original}$ 0.820 0.490 0.360 0.500 0.570 0.800 0.440 0.280 0.610 0.140 0.501
$\mathtt{SLD}$ 0.850 0.390 0.270 0.460 0.550 0.840 0.370 0.280 0.640 0.110 0.476
$\mathtt{AC}$ 0.350 0.230 0.220 0.250 0.220 0.620 0.120 0.150 0.210 0.110 0.248
$\mathtt{ESD}$ 0.020 0.020 0.090 0.160 0.000 0.060 0.030 0.040 0.060 0.010 0.049
$\mathtt{UCE}$ 0.030 0.060 0.080 0.070 0.030 0.020 0.040 0.020 0.020 0.040 0.041
$\mathtt{SA}$ 0.030 0.070 0.180 0.060 0.130 0.160 0.110 0.050 0.130 0.110 0.103
$\mathtt{RECELER}$ 0.000 0.010 0.050 0.030 0.010 0.000 0.020 0.020 0.010 0.030 0.018
$\mathtt{MACE}$ 0.010 0.110 0.090 0.050 0.010 0.070 0.050 0.050 0.090 0.040 0.057
Table 32: Attack robustness for each unlearning method across different target concepts. UnlearnDiffAtk (UDA) [zhang2024generate] is used.
$\mathtt{Celebrity}$
Angelina Jolie Ariana Grande Brad Pitt David Beckham Elon Musk Emma Watson Lady Gaga Leonardo DiCaprio Taylor Swift Tom Cruise Average
$\mathtt{Original}$ 0.902 0.839 0.842 0.821 0.883 0.924 0.753 0.863 0.926 0.851 0.860
$\mathtt{SLD}$ 0.022 0.067 0.021 0.124 0.074 0.034 0.097 0.133 0.080 0.042 0.069
$\mathtt{AC}$ 0.108 0.000 0.081 0.041 0.054 0.033 0.131 0.167 0.077 0.034 0.073
$\mathtt{ESD}$ 0.377 0.114 0.235 0.090 0.226 0.349 0.108 0.470 0.205 0.273 0.245
$\mathtt{UCE}$ 0.000 0.000 0.000 0.000 0.013 0.013 0.000 0.000 0.014 0.000 0.004
$\mathtt{SA}$ 0.000 0.000 0.000 0.000 0.000 0.000 0.013 0.011 0.000 0.000 0.002
$\mathtt{RECELER}$ 0.000 0.000 0.000 0.000 0.237 0.000 0.000 0.059 0.000 0.000 0.030
$\mathtt{MACE}$ 0.000 0.000 0.000 0.014 0.000 0.000 0.034 0.014 0.000 0.000 0.006
$\mathtt{Style}$
Andy Warhol Auguste Renoir Claude Monet Frida Kahlo Paul Cézanne Pablo Picasso Piet Mondrian Roy Lichtenstein Van Gogh Édouard Manet Average
$\mathtt{Original}$ 0.380 0.340 0.240 0.430 0.230 0.340 0.310 0.280 0.310 0.180 0.304
$\mathtt{SLD}$ 0.310 0.100 0.030 0.230 0.050 0.310 0.360 0.280 0.130 0.070 0.187
$\mathtt{AC}$ 0.320 0.140 0.080 0.390 0.060 0.190 0.230 0.210 0.100 0.160 0.188
$\mathtt{ESD}$ 0.250 0.010 0.010 0.080 0.020 0.130 0.220 0.110 0.010 0.010 0.085
$\mathtt{UCE}$ 0.270 0.120 0.060 0.230 0.090 0.260 0.200 0.110 0.110 0.060 0.151
$\mathtt{SA}$ 0.210 0.080 0.070 0.240 0.040 0.140 0.120 0.130 0.120 0.050 0.120
$\mathtt{RECELER}$ 0.100 0.000 0.000 0.020 0.000 0.030 0.040 0.000 0.010 0.010 0.021
$\mathtt{MACE}$ 0.270 0.050 0.070 0.030 0.010 0.160 0.130 0.090 0.060 0.040 0.091
$\mathtt{IP}$
Buzz Lightyear Homer Simpson Luigi Mario Mickey Mouse Pikachu Snoopy Sonic SpongeBob Stitch Average
$\mathtt{Original}$ 0.750 0.440 0.290 0.700 0.650 0.620 0.410 0.540 0.520 0.250 0.517
$\mathtt{SLD}$ 0.480 0.380 0.270 0.570 0.560 0.560 0.280 0.430 0.310 0.150 0.399
$\mathtt{AC}$ 0.290 0.350 0.280 0.490 0.480 0.490 0.250 0.460 0.180 0.190 0.346
$\mathtt{ESD}$ 0.080 0.080 0.070 0.190 0.120 0.110 0.020 0.170 0.030 0.010 0.088
$\mathtt{UCE}$ 0.020 0.000 0.040 0.060 0.000 0.070 0.020 0.010 0.020 0.000 0.024
$\mathtt{SA}$ 0.060 0.110 0.070 0.170 0.120 0.210 0.150 0.170 0.110 0.130 0.130
$\mathtt{RECELER}$ 0.010 0.020 0.000 0.040 0.010 0.010 0.010 0.000 0.000 0.010 0.011
$\mathtt{MACE}$ 0.030 0.040 0.090 0.070 0.210 0.100 0.020 0.100 0.020 0.060 0.074
Table 33: Attack robustness for each unlearning method across different target concepts. Unlearning or Concealment (UoC) [sharma2024unlearning] is used.